# Detecting Many-Body Scars from Fisher Zeros

### v2

@@ Abstract
The far-from-equilibrium dynamics of certain interacting quantum systems still defy precise understanding. One example is the so-called quantum many-body scars (QMBSs), where a set of energy eigenstates evade thermalization to give rise to long-lived oscillations. Despite the success of viewing scars from the perspectives of symmetry, commutant algebra and quasiparticles, it remains a challenge to elucidate the mechanism underlying all QMBS and to distinguish them from other forms of ergodicity breaking. In this work, we introduce an alternative route to detect and diagnose QMBS based on Fisher zeros, i.e., the patterns of zeros of the analytically continued partition function $Z$ on the complex $\beta$ (inverse temperature) plane. For systems with scars, a continuous line of Fisher zeros will appear off the imaginary $\beta$ axis and extend upward, separating the $\beta$-plane into regions with distinctive thermalization behaviors. This conjecture is motivated from interpreting the complex $Z$ as the return amplitude of the thermofield double state, and it is validated by analyzing two models with QMBS, the $\bar{P}X\bar{P}$ model and the Ising chain in external fields. These models also illustrate the key difference between QMBS and strong ergodicity breaking including their distinctive renormalization group flows on the complex $\beta$-plane. This “statistical mechanics” approach places QMBS within the same framework of thermal and dynamical phase transitions. It has the advantage of spotting scars without exhaustively examining each individual quantum state.
@@
Introduction. The Eigenstate Thermalization Hypothesis (ETH) postulates that an isolated interacting quantum system typically thermalizes [1, 2]. But there are known cases of ETH violation, even in non-integrable systems free of disorder. A prominent example is quantum many-body scars (QMBSs) [3, 4, 5, 6, 7] which have been observed on various platforms [8, 9, 10, 11]. Roughly speaking, a set of energy eigenstates, usually equally spaced to form a tower structure, become decoupled from the rest of the spectrum and give rise to persistent oscillations. The mechanism by which QMBS evade ETH has been attributed to Krylov-restricted Hilbert space fragmentation, spectrum-generating algebras, commutant algebras, or projector embeddings [12, 13, 14, 15], and they can also be understood from a quasiparticle perspective [16]. A unifying theoretical framework for all known QMBS, however, is still lacking.
@@
Key properties of QMBS are typically explored by examining the candidate Hamiltonian including its energy spectrum, eigenstates, the entanglement entropy, and Loschmidt echo. As a result, to find new instances of scar states, one must hand-select a few states and monitor their dynamics. The exhaustive search can turn prohibitively expensive for large system sizes, evocative of finding a needle in a haystack. Thus it is desirable to develop a diagnostic device that can spot the existence of QMBS while avoiding the inspection of each individual state. Compared to QMBS, strong breaking of ETH (SBETH) features more profound deviations from thermal equilibrium [17]. Examples include oscillations described by quasiparticles [18, 19, 20] and continuous time crystals [21, 22, 23, 24, 25, 26, 27] which spontaneously break time-translation symmetry. Possible connection between QMBS and SBETH has been suggested [28, 29]. But their coexistence and interplay remain poorly understood.
Figure 1: (Schematic) A synopsis of possible locations of Fisher zeros, the zeros of $Z=\mathrm{Tr}e^{-\beta H}$ on the complex $\beta=\beta_{r}+i\beta_{i}$ plane, in the thermodynamic limit. The thin line illustrates typical Fisher zeros in classical models, its intersection with the $\beta_{r}$-axis (the empty circle) marks the TPT point. Solid circles on the $\beta_{i}$-axis mark the DQPT points. The thick line touching the $\beta_{i}$-axis signals SBETH, while the line of zeros off the $\beta_{i}$-axis extending upward indicates QMBS.
@@
Motivated by these challenges, in this paper, we propose and demonstrate an alternative approach to QMBS. It is based on Fisher zeros originally invented to understand the singularities of equilibrium partition functions and thermal phase transitions (TPTs) [30, 31, 32]. The dynamical decoupling of scars from the rest of the spectrum is reminiscent of phase separation and hints that the tools from statistical mechanics may help yield fresh insights. To formalize this idea, we follow Fisher and analytically continue the partition function $Z=\sum_{n}e^{-\beta E_{n}}$ by allowing $\beta$ to take on complex values, $\beta=\beta_{r}+i\beta_{i}$. Then $Z$ becomes complex-valued, $Z=\sum_{n}e^{-\beta_{r}E_{n}}e^{-itE_{n}}$, which describes the time evolution of a thermal/mixed state at inverse temperature $\beta_{r}$ and at time $t=\beta_{i}$. The modulus square of $Z$, known as the spectral form factor, has been widely used in the study of quantum chaos [33]. To further connect $Z$ to pure state dynamics, we can purify the mixed state by enlarging the Hilbert space, e.g., by making a copy $R$ of the original system $L$ and defining the thermofield double (TFD) state $|\Psi(\beta_{r},0)\rangle=\sum_{n}e^{-\beta_{r}E_{n}/2}|n\rangle_{L}\otimes|n\rangle_{R}$ [34] (the thermal state is obtained by partial trace of $\Psi$ over subsystem $R$). Then $Z$ is nothing but the return amplitude of the pure TFD state evolving under the single Hamiltonian $H_{L}\otimes 1_{R}$, $Z(\beta_{r},t)\sim\langle\Psi(\beta_{r},0)|\Psi(\beta_{r},t)\rangle$, where the normalization factor is dropped for brevity and $\beta_{r}$ serves as a parameter of $\Psi$ that controls the weights of various eigenstates $|n\rangle$ [35]. The scar states contained in system $L$, if any, will continue to evade thermalization in the doubled system, for the added subsystem $R$ acts merely as a bath. Thus, the existence of scars will manifest in the return amplitude $Z(\beta_{r},t)$ at long times if the initial state $|\Psi(\beta_{r},0)\rangle$ has sufficient overlap (dictated by amplitudes $\propto e^{-\beta_{r}E_{n}/2}$) with the scar states. In other words, $Z$ cannot be a smooth single-valued function free of qualitative changes when $\beta_{r}$ is varied across the complex $\beta$-plane.
Figure 2: Fisher zeroes and real-time dynamics in the $\bar{P}X\bar{P}$ model: (a), (c), and (e) depict the Fisher zeros computed from HOTRG for $g=0.9$, 0.4, and 0, respectively ($L=64$). The blue and grey lines represent the solutions of $\mathrm{Re}Z=0$ and $\mathrm{Im}Z=0$. Their intersections (black dots) give the Fisher zeroes. (b), (d), and (f) show the frequency spectra of the AF state dynamics obtained from ED for system size $L=12$. The vertical dashed lines correspond to the first energy gap. The insets of each panel display the real-time oscillation curves of the AF (upper) and FM states (lower), with time in units of $2\pi$. (g) illustrates the two distinct regions of thermalization behavior: QMBS ($0\leq g\lesssim 0.3$) and SBETH ($10.3\lesssim g<$ 1).
@@
A conjecture regarding scars and Fisher zeros. These considerations lead to our main conjecture: for a system with QMBS, there exists at least a line of Fisher zeros that lies off the $\beta_{i}$-axis and extends to large $\beta_{i}$ (the long-time limit). As illustrated in Fig. 1, the line of zeros (labelled by QMBS) extends upward and marks the singularity of $f\equiv\ln|Z|$ in the limit of large system sizes. When the line is crossed, e.g., by increasing $\beta_{r}$ in the horizontal direction, the long-time dynamics of the TFD state will experience a qualitative change. Before we present numerical evidence to support this conjecture, it is illuminating to compare the location and implication of Fisher zeros in the present setting to their previous applications. In his seminar work on classical statistical models, Fisher observed that the zeros of $Z(\beta_{r},\beta_{i})$ pinching the $\beta_{r}$-axis marks a TPT. In recent years, the concept of the complex-valued partition function and its zeros has been extended to the field of quantum many-body physics. For example, we have shown that the vanishing of Fisher zero lines at infinite $\beta$ identifies the quantum critical point of the one-dimensional transverse-field Ising model [36, 37, 38]. In quench dynamics, following an initial state $\psi_{i}$, one can define a boundary partition function $Z_{i}(\beta_{r},\beta_{i})=\langle\psi_{i}|e^{-(\beta_{r}+i\beta_{i})H}|\psi_{i}\rangle$. The zeros of $Z_{i}$ (or the singularity of $f_{i}\equiv\ln|Z_{i}|$) on the $\beta_{i}$-axis then mark dynamical quantum phase transitions (DQPTs) [39, 40, 41, 42]. The empty and solid circles in Fig. 1 identify TPT and DQPT, respectively. In contrast, the zeros associated with QMBS are located away from the real and the imaginary $\beta$ axes. To our knowledge, the link between these Fisher zeros and breakdown of ETH has not been recognized before.
@@
We emphasize that the pattern of Fisher zeros on the complex $\beta$-plane is model dependent and may appear more complicated than the simple schematic in Fig. 1. In particular, it may contain additional features associated with SBETH [28] coexisting with QMBS. A typical scenario is shown in Fig. 1, where lines of Fisher zeros labeled by SBETH repeatedly intersect the $\beta_{i}$-axis, so $Z$ exhibits more pronounced non-analytical behaviors near the $\beta_{i}$-axis. To validate these conjectures, we first construct a generalized $PXP$ model to incorporate a crossover from QMBS to SBETH. Then, a nonintegrable model with Ising interactions is used to explore the rich interplays of the Fisher zeros when both QMBS and SBETH are present.
@@
The $\bar{P}X\bar{P}$ model. We generalize the well-known $PXP$ model [3, 43, 44], which describes atoms in the Rydberg-blockade regime, to the following $\bar{P}X\bar{P}$ model,
@@
We employ the tensor network method [45, 46, 47] to compute the complex $Z$ and its Fisher zeros. Specifically, using Trotter decomposition, the partition function of a one-dimensional quantum system of length $L$ can be mapped to a two-dimensional tensor network of size $L\times N$, where $N=\beta/\tau$ with the Trotter time $\tau$ [48]. In our calculations, we fix $N=1024$. We utilize the high order tensor renormalization group (HOTRG) algorithm [49] extended to complex $\beta$ [50] to contract this two-dimensional network and determine the locations where the real and imaginary parts of $Z$ vanish, thereby identifying the Fisher zeros through their intersections. The density of Fisher zeros increases with $L$ and usually form continuous lines in the thermodynamic limit [51]. To capture non-thermal oscillatory dynamics, we use exact diagonalization (ED) to perform long-time dynamical evolution $|\psi(\beta_{i})\rangle=\exp(-i\beta_{i}H)|\psi_{0}\rangle$ with a chosen initial state $\psi_{0}$, e.g., the antiferromagnetic (AF) or ferromagnetic (FM) state. Here $\beta_{i}$ goes up to $40\pi$ and plays the role of time $t$, and the system size $L=12$. We measure $\langle S_{z}\rangle=\langle\psi(\beta_{i})|S_{z}|\psi(\beta_{i})\rangle$ with $S_{z}=\sigma_{z}/2$ at any given site. The numerical results are shown in Fig. 2.
Figure 3: Key differences between SBETH and QMBS in the analytical structure of $Z$ and the Renormalization Group (RG) flow. Left: $f=\ln|Z|$, $f^{\prime}\equiv df/d\beta_{i}$, and $f_{i}=\ln|Z_{i}|$ for the AF (red) and FM (blue) states as functions of $\beta_{i}$ with $\beta_{r}=0$. In the SBETH case (a, $g=0.9$), $f$ becomes non-analytic over a broad interval of $\beta_{i}$. In the QMBS case (b, $g=0$), both $f$ and $f^{\prime}$ evolve smoothly and $f$ exibits only an inflection point. Right: the RG flows (arrows) for $g=0.9$ (c) and $g=0$ (d). The black lines connecting to the origin are the fixed-points defined by $S_{0}=1$. The contour lines are for different values of $S_{0}$ at $L=64$.
@@
At $g=1$, the Fisher zeros are located at isolated points $\beta_{i}=\pi/2+n\pi$ on the $\beta_{i}$-axis. Correspondingly, all states, including AF and FM, undergo Rabi oscillations. As $g$ is decreased, the interaction effects kick in and the Fisher zeros extend from the $\beta_{i}$-axis into the complex $\beta$-plane to form small line segments. The solutions of $\mathrm{Re}Z=0$ and $\mathrm{Im}Z=0$ remain dense near the $\beta_{i}$-axis [Fig. 2(a)]. Both AF and FM states exhibit oscillations, with frequencies corresponding to the system’s low-energy excitations. In the small $\beta_{r}$ region near the $\beta_{i}$-axis, a vertical path will cut through these lines of zeros, suggesting that thermalization is hindered at long times. The evasion of thermalization is closely related to the long-lived low-energy excitations. Therefore, this parameter regime can be identified as being dominated by SBETH [Fig. 2(b)]. As $g$ further decreases, additional line segments of Fisher zeros emerge, while the density of the solutions of $\mathrm{Re}Z=0$ and $\mathrm{Im}Z=0$ gradually decreases [Fig. 2(c)]. Accordingly, the main oscillation frequency of the AF state shifts away from the low-energy excitations, and a secondary peak emerges at roughly twice the frequency, reminiscent of the tower structure of scars. In contrast, the FM state ceases to oscillate with only weak spin-flip effects retained. This regime can be viewed as the crossover region between SBETH and QMBS [Fig. 2(d)]. Around $g\approx 0.3$, the Fisher zero lines begin to detach from the $\beta_{i}$-axis and connect vertically [Fig. 2(e)]. Dynamically, only the AF state maintains non-thermal oscillations, so the parameter regime $g\lesssim 0.3$ is dominated by QMBS [Fig. 2(f)]. Figure. 2(g) summarizes the crossover from QMBS to SBETH as $g$ is varied, with each region depicted by its distinct structures of Fisher zeros.
Mountain range and water basin. To gain a deeper understanding of the location of Fisher zeros for QMBS and its qualitative difference from SBETH, we examine the analytical properties of $Z$ along the $\beta_{i}$-axis. Recall $Z=\sum_{i}Z_{i}$ where $\{|\psi_{i}\rangle\}$ form a complete basis and the zeros of $Z_{i}$ (DQPT) mark the singularities of $f_{i}=\ln|Z_{i}|$. In the case of SBETH [Fig. 3(a)], $f$ becomes non-analytic over finite intervals (shaded regions), due to the collective effect of a broad set of states developing close-packed singularities in $f_{i}$. Interestingly, one of the singular points of the AF state (red curve) is very close to the boundary of the non-analytic region. In contrast, for the QMBS case both $f$ and $f^{\prime}$ remain smooth [Fig. 3(b)], different $Z_{i}$ exhibit different patterns of zeros in the complex $\beta$-plane [51], and the dynamics of $Z_{i}(\beta_{r}=0,\beta_{i})$ is also different for scar and non-scar states. In this case, the singularities of the non-scar FM state have little effect on $f$, while a singular point of the scar AF state coincides with the inflection point of $f$ (shaded narrow region). The inflection point also corresponds to the projection of the lowest Fisher zero (see Fig. 2e) onto the $\beta_{i}$-axis. This “shadow effect” is analogous to how the rightmost Fisher zeros near the $\beta_{r}$-axis influence thermodynamic quantities such as specific heat [38].
The distinction between strong and weak ergodicity breaking can also be elucidated from renormalization group (RG). It has been noted that Fisher zeros form the boundary of RG flows on the complex $\beta$-plane [52, 53]. As shown in  [51], the “equipotential line” defined by $S_{0}=|Z(\beta_{r},t)/Z(0,0)|^{2}=1$ converges to a continuous line in $\beta$-plane as the system size $L\rightarrow\infty$, providing a convenient “high-temperature” (since it is connected to the origin $\beta=0$) fixed line for RG. The general trend of RG flows in Fig. 3 can be appreciated by a geological analogy: water flows toward the $S_{0}=1$ line (river in water basin) from near the Fisher zeros (mountain range). For SBETH, the Fisher zeros cross the $\beta_{i}$-axis, allowing the RG flow to reach the $S_{0}=1$ line throughout the $\beta$-plane [Fig. 3(c)]. The periodic return along the time ($\beta_{i}$) axis clearly violates the ETH. In the QMBS case, there exists a large plateau region between the zero line and the $\beta_{i}$-axis where the RG flows (not shown) are blocked by the “mountain range” of Fisher zeros and cannot reach the $S_{0}=1$ fixed line [Fig. 3(d)]. Thus the “mountain range” of Fisher zeros divides the $\beta$-plane into regions with distinctive long-time behaviors in $Z$, in agreement with the intuitive argument used earlier to motivate our conjecture.
@@
Ising chain with external fields. To demonstrate the rich interplay between QMBS and SBETH and the utility of diagnosing QMBS through Fisher zeros, we examine another many-body Hamiltonian, the one-dimensional Ising chain in external fields with PBC,
@@
[…]
@@
QMBS emerges when both $h_{l}$ and $g_{t}$ are present. The exact locations of Fisher zeros can be found for $h_{l}=0$ [36]. For instance, for $g_{t}<1$, the open lines and closed loops of zeros alternately appear in the $\beta$-plane. When $h_{l}$ is introduced, the closed loops break open, setting free the meson quasiparticles previously confined within the loops [36]. This is shown in Fig. 4(c) for the example of $g_{t}=0.5$ and $h_{l}=0.1$. Correspondingly, a set of Fisher zeros are displaced away from the $\beta_{i}$-axis by $h_{l}$ to form a continuous line (black dots). This is the hallmark of QMBS, and rather similar to the $\bar{P}X\bar{P}$ model near $g=0$. The coexistence of these two types of Fisher zeros, namely the broken loops and the emergent line extending to large $\beta_{i}$, suggests that in strongly correlated models, quasiparticles can influence scar phenomena. We choose the FM state, which does not undergo DQPT, as an example of the time evolution. Its oscillatory behavior and the corresponding frequencies are shown in Fig. 4(d). We find that the peak at frequency 2.9, marked by a red arrow, remains unchanged with system size and away from the system’s lowest-energy excitation. This peak corresponds exactly to the scar effect induced by meson excitations [6]. Although other nearly equally spaced peaks arise due to finite-size effects, they build upon the meson peak [51].
@@
Generalization and outlook. We have confirmed our conjecture which relates Fisher zeros to QMBS using two models, Eqs. (1) and (2). We have also identified the distinct characteristics of weak and strong ergodicity breaking: for QMBS the Fisher zeros are off the $\beta_{i}$-axis extending to large $\beta_{i}$, while for SBETH the Fisher zeros intersect the $\beta_{i}$-axis. This approach can be applied to other models. One example is provided in [51], where we solve a cluster spin model closely related to 1+1D lattice gauge theory [54] that exhibits QMBS [55, 56]. Its Fisher zero configurations again support the validity of our conjecture. These three model studies offer strong evidence that the unifying link between Fisher zeros and thermalization breakdown is rather general.

### v3

@@ Abstract
The far-from-equilibrium dynamics of certain interacting quantum systems still defy precise understanding. One example is the so-called quantum many-body scars (QMBSs), where a set of energy eigenstates evade thermalization to give rise to long-lived oscillations. Despite the success of viewing scars from the perspectives of symmetry, commutant algebra, and quasiparticles, it remains a challenge to elucidate the mechanism underlying all QMBS and to distinguish them from other forms of ergodicity breaking. In this work, we introduce an alternative route to detect and diagnose QMBS based on Fisher zeros, i.e., the patterns of zeros of the analytically continued partition function $Z$ on the complex $\beta$ (inverse temperature) plane. For systems with scars, a continuous line of Fisher zeros will appear off the imaginary $\beta$ axis and extend upward, separating the $\beta$ plane into regions with distinctive thermalization behaviors. This conjecture is motivated from interpreting the complex $Z$ as the return amplitude of the thermofield double state, and it is validated by analyzing two models with QMBS, the $\bar{P}X\bar{P}$ model and the Ising chain in external fields. These models also illustrate the key difference between QMBS and strong ergodicity breaking including their distinctive renormalization group flows on the complex $\beta$ plane. This “statistical mechanics” approach places QMBS within the same framework of thermal and dynamical phase transitions. It has the advantage of spotting scars without exhaustively examining each individual quantum state.
@@
Introduction—The eigenstate thermalization hypothesis (ETH) postulates that an isolated interacting quantum system typically thermalizes [1, 2]. But there are known cases of ETH violation, even in nonintegrable systems free of disorder. A prominent example is quantum many-body scars (QMBSs) [3, 4, 5, 6, 7], which have been observed on various platforms [8, 9, 10, 11]. Roughly speaking, a set of energy eigenstates, usually equally spaced to form a tower structure, become decoupled from the rest of the spectrum and give rise to persistent oscillations. The mechanism by which QMBS evade ETH has been attributed to Krylov-restricted Hilbert space fragmentation, spectrum-generating algebras, commutant algebras, or projector embeddings [12, 13, 14, 15], and they can also be understood from a quasiparticle perspective [16]. A unifying theoretical framework for all known QMBS, however, is still lacking.
@@
Key properties of QMBS are typically explored by examining the candidate Hamiltonian including its energy spectrum, eigenstates, the entanglement entropy, and Loschmidt echo. As a result, to find new instances of scar states, one must hand select a few states and monitor their dynamics. The exhaustive search can turn prohibitively expensive for large system sizes, evocative of finding a needle in a haystack. Thus it is desirable to develop a diagnostic device that can spot the existence of QMBS while avoiding the inspection of each individual state. Compared to QMBS, strong breaking of ETH (SBETH) features more profound deviations from thermal equilibrium [17]. Examples include oscillations described by quasiparticles [18, 19, 20] and continuous time crystals [21, 22, 23, 24, 25, 26, 27], which spontaneously break time-translation symmetry. Possible connection between QMBS and SBETH has been suggested [28, 29]. But their coexistence and interplay remain poorly understood.
@@
Motivated by these challenges, in this Letter, we propose and demonstrate an alternative approach to QMBS. It is based on Fisher zeros originally invented to understand the singularities of equilibrium partition functions and thermal phase transitions (TPTs) [30, 31, 32]. The dynamical decoupling of scars from the rest of the spectrum is reminiscent of phase separation and hints that the tools from statistical mechanics may help yield fresh insights. To formalize this idea, we follow Fisher and analytically continue the partition function $Z=\sum_{n}e^{-\beta E_{n}}$ by allowing $\beta$ to take on complex values, $\beta=\beta_{r}+i\beta_{i}$. Then $Z$ becomes complex valued, $Z=\sum_{n}e^{-\beta_{r}E_{n}}e^{-itE_{n}}$, which describes the time evolution of a thermal or mixed state at inverse temperature $\beta_{r}$ and at time $t=\beta_{i}$. The modulus square of $Z$, known as the spectral form factor, has been widely used in the study of quantum chaos [33]. To further connect $Z$ to pure state dynamics, we can purify the mixed state by enlarging the Hilbert space, e.g., by making a copy $R$ of the original system $L$ and defining the thermofield double (TFD) state $|\Psi(\beta_{r},0)\rangle=\sum_{n}e^{-\beta_{r}E_{n}/2}|n\rangle_{L}\otimes|n\rangle_{R}$ [34] (the thermal state is obtained by partial trace of $\Psi$ over subsystem $R$). Then $Z$ is nothing but the return amplitude of the pure TFD state evolving under the single Hamiltonian $H_{L}\otimes 1_{R}$, $Z(\beta_{r},t)\sim\langle\Psi(\beta_{r},0)|\Psi(\beta_{r},t)\rangle$, where the normalization factor is dropped for brevity and $\beta_{r}$ serves as a parameter of $\Psi$ that controls the weights of various eigenstates $|n\rangle$ [35].
The scar states contained in system $L$, if any, will continue to evade thermalization in the doubled system, for the added subsystem $R$ acts merely as a bath. Thus, the existence of scars will manifest in the return amplitude $Z(\beta_{r},t)$ at long times if the initial state $|\Psi(\beta_{r},0)\rangle$ has sufficient overlap (dictated by amplitudes $\propto e^{-\beta_{r}E_{n}/2}$) with the scar states. In other words, $Z$ cannot be a smooth single-valued function free of qualitative changes when $\beta_{r}$ is varied across the complex $\beta$ plane.
Figure 1: (Schematic) A synopsis of possible locations of Fisher zeros, the zeros of $Z=\mathrm{Tr}e^{-\beta H}$ on the complex $\beta=\beta_{r}+i\beta_{i}$ plane in the thermodynamic limit. The thin line illustrates typical Fisher zeros in classical models, its intersection with the $\beta_{r}$ axis (the empty circle) marks the TPT point. Solid circles on the $\beta_{i}$ axis mark the DQPT points. The dashed line illustrates one scenario of SBETH with the Fisher zeros touching the $\beta_{i}$ axis. The thick line of zeros off the $\beta_{i}$ axis extending upward indicates QMBS. Darker regions feature larger thermodynamic and smaller quantum fluctuations.
@@
A conjecture regarding scars and Fisher zeros—These considerations lead to our main conjecture: for a system with QMBS, there exists at least a line of Fisher zeros that lies off the $\beta_{i}$ axis and extends to large $\beta_{i}$ (the long-time limit). As illustrated in Fig. 1, the line of zeros (labeled by QMBS) extends upward and marks the singularity of $f\equiv\ln|Z|$ in the limit of large system sizes. When the line is crossed, e.g., by increasing $\beta_{r}$ in the horizontal direction, the long-time dynamics of the TFD state will experience a qualitative change. Before we present numerical evidence to support this conjecture, it is illuminating to compare the location and implication of Fisher zeros in the present setting to their previous applications. In his seminar work on classical statistical models, Fisher observed that the zeros of $Z(\beta_{r},\beta_{i})$ pinching the $\beta_{r}$ axis marks a TPT. In recent years, the concept of the complex-valued partition function and its zeros has been extended to the field of quantum many-body physics. For example, we have shown that the vanishing of Fisher zero lines at infinite $\beta$ identifies the quantum critical point of the one-dimensional transverse-field Ising model [36, 37, 38]. In quench dynamics, following an initial state $\psi_{i}$, one can define a boundary partition function $Z_{i}(\beta_{r},\beta_{i})=\langle\psi_{i}|e^{-(\beta_{r}+i\beta_{i})H}|\psi_{i}\rangle$. The zeros of $Z_{i}$ (or the singularity of $f_{i}\equiv\ln|Z_{i}|$) on the $\beta_{i}$ axis then mark dynamical quantum phase transitions (DQPTs) [39, 40, 41, 42]. The empty and solid circles in Fig. 1 identify TPT and DQPT, respectively. In contrast, the zeros associated with QMBS are located away from the real and the imaginary $\beta$ axes. To our knowledge, the link between these Fisher zeros and breakdown of ETH has not been recognized before.
@@
We emphasize that the pattern of Fisher zeros on the complex $\beta$ plane is model dependent and may appear more complicated than the simple schematic in Fig. 1. In particular, it may contain additional features associated with SBETH [28] coexisting with QMBS. A typical scenario is shown in Fig. 1, where lines of Fisher zeros labeled by SBETH repeatedly intersect the $\beta_{i}$ axis, so $Z$ exhibits more pronounced non-analytical behaviors near the $\beta_{i}$ axis. To validate the conjecture regarding QMBS, we first construct a generalized $PXP$ model to incorporate a crossover from QMBS to SBETH. Then, a nonintegrable model with Ising interactions is used to explore the rich interplays of the Fisher zeros when both QMBS and SBETH are present.
@@
The $\bar{P}X\bar{P}$ model—We generalize $PXP$-like models [3, 43, 44, 45, 46, 47], which describe atoms in the Rydberg-blockade regime, to the following $\bar{P}X\bar{P}$ model,
@@
Figure 2: Fisher zeroes and real-time dynamics in the $\bar{P}X\bar{P}$ model: (a), (c), and (e) depict the Fisher zeros computed from HOTRG for $g=0.9$, 0.4, and 0, respectively ($L=64$). The blue and gray lines represent the solutions of $\mathrm{Re}Z=0$ and $\mathrm{Im}Z=0$. Their intersections (black dots) give the Fisher zeroes. (b), (d), and (f) show the frequency spectra of the AFM state dynamics obtained from ED for system size $L=12$. The vertical dashed lines correspond to the first energy gap. The insets of each panel display the real-time oscillation curves of the AFM (upper) and FM states (lower), with time in units of $2\pi$. (g) The two distinct regions of thermalization behavior: QMBS ($0\leq g\lesssim 0.3$) and SBETH ($0.3\lesssim g<1$).
@@
We employ the tensor network method [48, 49, 50] to compute the complex $Z$ and its Fisher zeros. Specifically, using Trotter decomposition, the partition function of a one-dimensional quantum system of length $L$ can be mapped to a two-dimensional tensor network of size $L\times N$, where $N=\beta/\tau$ with the Trotter time $\tau$ [51]. In our calculations, we fix $N=1024$. We utilize the high order tensor renormalization group (HOTRG) algorithm [52] extended to complex $\beta$ [53] to contract this two-dimensional network and determine the locations where the real and imaginary parts of $Z$ vanish, thereby identifying the Fisher zeros through their intersections. The density of Fisher zeros increases with $L$ and usually form continuous lines in the thermodynamic limit [54]. To capture nonthermal oscillatory dynamics, we use exact diagonalization (ED) to perform long-time dynamical evolution $|\psi(\beta_{i})\rangle=\exp(-i\beta_{i}H)|\psi_{0}\rangle$ with a chosen initial state $\psi_{0}$, e.g., the antiferromagnetic (AFM) or ferromagnetic (FM) state. Here $\beta_{i}$ goes up to $40\pi$ and plays the role of time $t$, and the system size $L=12$. We measure $\langle S_{z}\rangle=\langle\psi(\beta_{i})|S_{z}|\psi(\beta_{i})\rangle$ with $S_{z}=\sigma_{z}/2$ at any given site. The numerical results are shown in Fig. 2.
@@
At $g=1$, the Fisher zeros are located at isolated points $\beta_{i}=\pi/2+n\pi$ on the $\beta_{i}$-axis. Correspondingly, all states, including AFM and FM, undergo Rabi oscillations. As $g$ is decreased, the interaction effects kick in and the Fisher zeros extend from the $\beta_{i}$ axis into the complex $\beta$ plane to form small line segments. The solutions of $\mathrm{Re}Z=0$ and $\mathrm{Im}Z=0$ remain dense near the $\beta_{i}$ axis [Fig. 2(a)]. Both AFM and FM states exhibit oscillations, with frequencies corresponding to the system’s low-energy excitations. In the small $\beta_{r}$ region near the $\beta_{i}$ axis, a vertical path will cut through these lines of zeros, suggesting that thermalization is hindered at long times. The evasion of thermalization is closely related to the long-lived low-energy excitations. Therefore, this parameter regime can be identified as being dominated by SBETH [Fig. 2(b)]. As $g$ further decreases, additional line segments of Fisher zeros emerge, while the density of the solutions of $\mathrm{Re}Z=0$ and $\mathrm{Im}Z=0$ gradually decreases [Fig. 2(c)]. Accordingly, the main oscillation frequency of the AFM state shifts away from the low-energy excitations, and a secondary peak emerges at roughly twice the frequency, reminiscent of the tower structure of scars. In contrast, the FM state ceases to oscillate with only weak spin-flip effects retained. This regime can be viewed as the crossover region between SBETH and QMBS [Fig. 2(d)]. Around $g\approx 0.3$, the Fisher zero lines begin to detach from the $\beta_{i}$ axis and connect vertically [Fig. 2(e)]. Dynamically, only the AFM state maintains nonthermal oscillations, so the parameter regime $g\lesssim 0.3$ is dominated by QMBS [Fig. 2(f)]. Figure. 2(g) summarizes the crossover from QMBS to SBETH as $g$ is varied, with each region depicted by its distinct structures of Fisher zeros. More theoretical arguments for why Fisher zeros are a general diagnostic tool for QMBS are presented in the End Matter.
@@
Ising chain with external fields—To demonstrate the rich interplay between QMBS and SBETH and the utility of diagnosing QMBS through Fisher zeros, we examine another many-body Hamiltonian, the one-dimensional Ising chain in external fields with PBC,
@@
We first consider the $g_{t}=0$ case with only the longitudinal field present. In this limit $Z$ and Fisher zeros can be determined exactly. When $h_{l}=0$, the Fisher zeros form lines parallel to the $\beta_{r}$ axis, extending to $\beta_{r}=+\infty$. Once $h_{l}$ is introduced, these zeros terminate at finite values of $\beta_{r}$, yielding a structure similar to that observed near $g=1$ in the $\bar{P}X\bar{P}$ model. Figure 3(a) shows the configuration of Fisher zeros (black dots) for the case $h_{l}=0.2$: in the complex $\beta$ plane, they cross the $\beta_{i}$ axis in periodic fashion. The termination points of the zeros from different line sectors form an envelope curve with a larger period, indicating a proportional relationship between the two periodicities. We select $|\cdots\rightarrow\cdots\rangle$, i.e., the direct product of all the $\sigma^{x}_{j}$ eigenstates, and illustrate its dynamical evolution (inset) and the corresponding frequency spectrum in Fig. 3(b). The intricate SBETH oscillations have frequencies matching certain excitations (indicated by dash lines) in the system’s spectrum. These frequency modes also align with the periodicities of the Fisher zeros and the envelope curve in Fig. 3(a).
Figure 3: Fisher zeros and real-time dynamics of the Ising chain Eq. (2). (a) The Fisher zeros (black dots) for $g_{t}=0$, $h_{l}=0.2$ ($L=32$). (b) The frequency spectrum of the $|\cdots\rightarrow\cdots\rangle$ state for $L=12$. Dashed lines correspond to the energy levels in the spectrum. (c) The Fisher zeros of the case $g_{t}=0.5$, $h_{l}=0.1$, and $L=128$. Only the zeros describing the scars are marked by black dots. (d) The frequency spectrum of the FM state. Dashed lines correspond to the first excited state. The red arrow points to the scar peak, while other peaks arising from finite-size effects form a tower structure.
@@
QMBS emerges when both $h_{l}$ and $g_{t}$ are present. The exact locations of Fisher zeros can be found for $h_{l}=0$ [36]. For instance, for $g_{t}<1$, the open lines and closed loops of zeros alternately appear in the $\beta$ plane. When $h_{l}$ is introduced, the closed loops break open, setting free the meson quasiparticles previously confined within the loops [36]. This is shown in Fig. 3(c) for the example of $g_{t}=0.5$ and $h_{l}=0.1$. Correspondingly, a set of Fisher zeros are displaced away from the $\beta_{i}$ axis by $h_{l}$ to form a continuous line (black dots). This is the hallmark of QMBS, and rather similar to the $\bar{P}X\bar{P}$ model near $g=0$. The coexistence of these two types of Fisher zeros, namely, the broken loops and the emergent line extending to large $\beta_{i}$, suggests that in strongly correlated models, quasiparticles can influence scar phenomena. We choose the FM state, which does not undergo DQPT, as an example of the time evolution. Its oscillatory behavior and the corresponding frequencies are shown in Fig. 3(d). We find that the peak at frequency 2.9, marked by a red arrow, remains unchanged with system size and away from the system’s lowest-energy excitation. This peak corresponds exactly to the scar effect induced by meson excitations [6]. Although other nearly equally spaced peaks arise due to finite-size effects, they build upon the meson peak [54].
@@
Generalization and outlook—We have confirmed our conjecture which relates Fisher zeros to QMBS using two models, Eqs. (1) and (2). We have also identified the distinct characteristics of weak and strong ergodicity breaking: for QMBS the Fisher zeros are off the $\beta_{i}$ axis extending to large $\beta_{i}$, while for SBETH the Fisher zeros intersect the $\beta_{i}$ axis. This approach can be applied to other models. One example is provided in [54], where we solve a cluster spin model closely related to 1+1D lattice gauge theory [55] that exhibits QMBS [56, 57]. Its Fisher zero configurations again support the validity of our conjecture. These three model studies offer strong evidence that the unifying link between Fisher zeros and thermalization breakdown is rather general.